T=2π√k²+h²/gh

Where h=5 and T=1.270

Find g



1
Expert's answer
2021-03-23T11:12:43-0400
"T=2\pi\sqrt{\frac{l}{g}}\\1.27=2\pi\sqrt{\frac{5}{g}}\\g=122\frac{m}{s^2}"